OpenStudy (anonymous):

Find the exact value logv3 5*logv5 81

OpenStudy (campbell_st):

just use change of base \[\log_{a}(b) = \frac{\ln(b)}{\ln(a)}\] change so they are both base 5 or both base 3

OpenStudy (campbell_st):

I think its easier if you change the 2nd log to base 3 ... it falls out easier that way